Hubble Repairs Working Well

Hubble Status Report #2

NASA Hubble Space Telescope Status Update #1

"The Hubble Space Telescope team completed switching the required hardware modules to their B-sides about 9:30 a.m. this morning and received telemetry that verified they had good data. Everything at this point looks good."

NASA Hubble Space Telescope Daily Status Report #4715

"At 284/14:25:46 UTC, the SSM HV Protection capability through the NSSC-1 was successfully restored via Ops Request 18307-0. Executive RTCS 21 was enabled to send notification of an HV Protect event through the setting of the former Mechanism Motion flag, and NSSC-1 Global Event Flag 12 was set to allow this transfer across the PIT to the SSM."

  • submit to reddit


The 2020 AAS Goddard Symposium
Small Satellites Conference 2020
Humans to Mars Summit 2020

Join our mailing list

Commercialization: Monthly Archives

Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on October 16, 2008 10:09 AM.

President Signs NASA Authorization Act of 2008 was the previous entry in this blog.

NASA Advisory Council Has A New Chairman is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.